Find all solutions to the equation cos x = -1.
Answer by stanbon(48564)

The pricipal solution is x = 180 degrees.
You cannot list all the solutions unless you like the following:
x = 180 + 360n where n= 1,2,3,..... or n=-1,-2,-3,....
Cheers,
Stan H.